OpenStudy (anonymous):

Define a system of equations. Solve systems of equations with two and three variables. During the 1998-1999 Little League season, the Tigers played 52 games. They won 12 more games than they lost. How many games did they win that season?

OpenStudy (dumbcow):

w+L = 52 w = L+12

OpenStudy (dumbcow):

(L+12) +L =52 2L +12 = 52 2L = 40 L = 20 W = 32

OpenStudy (anonymous):

Now I understand thank you.